# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ 6.2 Interim Allocation Probabilities Do Not Guarantee Robustness
In this section, we examine if the upper-bound restrictions on $p_{j}^{\{i,j\}}$ of Lemma 6.1 are necessary for the robustness guarantee or just specifying the interim allocations is enough. We will show that these restrictions are indeed necessary: we construct an example where the allocation probabilities $(p_{i}^{S})$ induce the desired interim probabilities of Theorem 4.2 but some agent $i$ cannot guarantee more than an $2\alpha_{i}$ fraction of her ideal utility robustly.
@@
$p_{i}^{S}=\begin{cases}1/|S|&\text{if $1\notin S$}\\ 1&\text{if $|S|\neq 2$ and $i=1$}\\ 0&\text{if $|S|\neq 2$ and $1\in S$ and $i\neq 1$}\\ 1/n&\text{if $|S|=2$ and $i=1$}\\ 1-1/n&\text{if $|S|=2$ and $i\neq 1$}\end{cases}.$ (7)
In words, we set the allocation probabilities such that if agent $1$ does not bid, we allocate uniformly at random, and if agent $1$ bids and more than one other agent bids, we allocate to agent $1$ with probability $1$, and if agent $1$ bids and exactly one other agent bids, we allocate to agent $1$ with probability $1/n$. We carefully picked these allocation probabilities such that they induce the same interim allocation probabilities $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ as in Theorem 4.2. In addition, the probabilities violate the restriction of Lemma 6.1 as much as possible: $p_{j}^{\{1,j\}}=1-1/n\gg 1/2$ when $n$ is large.
Even though the $p_{i}^{S}$ induce the correct interim allocation probabilities, they do not guarantee robustness by the following proposition, proved formally in Appendix D.
@@
Using the allocation probabilities $p_{i}^{S}$ as defined by (7), player $1$ does not have a $\lambda$-robust strategy for any $\lambda>\frac{n-1}{n^{2}}+\frac{1}{n}-O\left(\sqrt{\nicefrac{{\log T}}{{T}}}\right)$.
Proof sketch.
Suppose other players $i\neq 1$ take turns bidding, exactly one agent $i\neq 1$ bidding per round. Since they have total budget a $\frac{n-1}{n}$ fraction of the rounds, they can do this for all rounds but $T/n$. By how the $p_{i}^{S}$’s are set, agent $1$ can only win the item with probability $1/n$ on these rounds. This results in agent $1$’s utility to be extremely low. ∎
@@ 6.3 A $1/2$ Hardness Result
In this section, we show that our analysis for the robustness bound of Theorem 4.3 is approximately tight if the fair shares are small. Specifically, we show that the BRB mechanism cannot guarantee every agent more than a 1/$21/2+\sum_{i}\alpha_{i}^{2}/$ 2 fraction of her ideal utility. We show this for any allocation probabilities $p_{i}^{S}$ that the mechanism could use.
The following lemma stems from the strategy used by agents $j\neq i$ bidding in an anti-correlated fashion. They make sure that at most $1$ agent $j\neq i$ bids at each time by coordinating with each other. Each agent $j\neq i$ can bid for an $\alpha_{j}$ fraction of the time. Agents $j\neq i$ will not bid for a $1-\sum_{j\neq i}\alpha_{j}=\alpha_{i}$ fraction of the time, so agent $i$ can win these rounds with no competition. At each round that agent $j$ bids, agent $i$ can only win with probability $p_{i}^{\{i,j\}}$ conditioned on agent $i$ bidding. Summing the probabilities that agent $i$ can win conditioned on bidding over all times, we obtain the $\alpha_{i}+\sum_{j\neq i}\alpha_{j}p_{i}^{\{i,j\}}$ term below.
@@
When the slack parameters are set as $\delta_{k}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{k}T}}}$, no matter what value distributions agent $i$ has, if agent $i$has a $\lambda$-robust strategy has a $\lambda_{i}$-robust strategy, then
@@
We formally prove the lemma in Appendix D. We can use the bound in the above lemma to show that there is always an agent that cannot have a $\lambda$-robust strategy for $\lambda$ not much higher than $1/2$.
@@
When the slack parameters are set as $\delta_{k}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\beta_{k}T}}}$, no matter what value distributions the agents have, if every agent $i$ has a $\lambda$-robust strategy, then
@@
The theorem can be proved using Lemma 6.5 by summing the terms of that lemma. The full proof is included in Appendix D.
[…]
Appendix A Better Equilibrium Guarantees

### v2

@@ 5.2 Interim Allocation Probabilities Do Not Guarantee Robustness
In this section, we examine if the upper-bound restrictions on $p_{j}^{\{i,j\}}$ of Lemma˜5.1 are necessary for the robustness guarantee or just specifying the interim allocations is enough. We will show that these restrictions are indeed necessary: we construct an example where the allocation probabilities $(p_{i}^{S})$ induce the desired interim probabilities of Theorem˜3.5 but some agent $i$ cannot guarantee more than an $2\alpha_{i}$ fraction of her ideal utility robustly.
@@
[…]
In words, we set the allocation probabilities such that if agent $1$ does not bid, we allocate uniformly at random, and if agent $1$ bids and more than one other agent bids, we allocate to agent $1$ with probability $1$, and if agent $1$ bids and exactly one other agent bids, we allocate to agent $1$ with probability $1/n$. We carefully picked these allocation probabilities such that they induce the same interim allocation probabilities $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ as in Theorem˜3.5. In addition, the probabilities violate the restriction of Lemma˜5.1 as much as possible: $p_{j}^{\{1,j\}}=1-1/n\gg 1/2$ when $n$ is large.
Even though the $p_{i}^{S}$ induce the correct interim allocation probabilities, they do not guarantee robustness by the following proposition, proved formally in subsection˜E.7.
@@
Using the allocation probabilities $p_{i}^{S}$ as defined by (5.4), player $1$ does not have a $\lambda$-robust strategy for any $\lambda>\frac{n-1}{n^{2}}+\frac{1}{n}-O\left(\sqrt{\nicefrac{{\log T}}{{T}}}\right)$.
Proof 3 (Proof sketch).
Suppose other players $i\neq 1$ take turns bidding, exactly one agent $i\neq 1$ bidding per round. Since they have total budget a $\frac{n-1}{n}$ fraction of the rounds, they can do this for all rounds but $T/n$. By how the $p_{i}^{S}$’s are set, agent $1$ can only win the item with probability $1/n$ on these rounds. This results in agent $1$’s utility to be extremely low.
@@ 5.3 A $1/2$ Hardness Result
In this section, we show that our analysis for the robustness bound of Theorem˜3.6 is approximately tight if the fair shares are small. Specifically, we show that the BRB mechanism cannot guarantee every agent more than a $1/2+\sum_{i}\alpha_{i}^{2}/2$ fraction of her ideal utility. We show this for any allocation probabilities $p_{i}^{S}$ that the mechanism could use, not just the ones that induce optimal performance at equilibrium.
The following lemma is a generalization of Proposition˜5.4 and its intuition is the same as the proof sketch of Proposition˜5.4: it stems from the case where the other players $i\neq 1$ take turns bidding, exactly one agent $i\neq i$ bidding per round. Each agent $j\neq i$ can bid for an $\alpha_{j}$ fraction of the time. Agents $j\neq i$ will not bid for a $1-\sum_{j\neq i}\alpha_{j}=\alpha_{i}$ fraction of the time, so agent $i$ can win these rounds with no competition. At each round that agent $j$ bids, agent $i$ can only win with probability $p_{i}^{\{i,j\}}$ conditioned on agent $i$ bidding. Summing the probabilities that agent $i$ can win conditioned on bidding over all times, we obtain the $\alpha_{i}+\sum_{j\neq i}\alpha_{j}p_{i}^{\{i,j\}}$ term below.
@@
In the BRB mechanism, when the slack parameters are set as $\delta_{k}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{k}T}}}$, no matter what value distribution agent $i$ has, any $\lambda_{i}$-robust strategy of agent $i$ satisfies
@@
We formally prove the lemma in subsection˜E.7. We can use the bound in the above lemma to show that there is always an agent that cannot have a $\lambda$-robust strategy for a constant $\lambda$ greater than $1/2$ not depending on the fair shares $(\alpha_{i})$ or the time horizon $T$.
@@
In the BRB mechanism, when the slack parameters are set as $\delta_{k}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\beta_{k}T}}}$, no matter what value distributions the agents have, if every agent $i$ has a $\lambda$-robust strategy, then
@@
The theorem can be proved using Lemma˜5.5 by summing the terms of that lemma. The full proof is included in subsection˜E.8.
6 Computationally Efficient BRB Mechanism
In previous sections, we gave constraints on the allocation probabilities $p_{i}^{S}$ to obtain optimal equilibrium utility and $1/2$-robustness guarantees. Specifically, we showed that if we can make the interim allocations $p_{i}$ equal to $1-\prod_{j\in[n]}(1-\alpha_{j})$, i.e., if $\mu$ is the distribution of subsets of $[n]$ where each agent $i$ independently requests with probability $\alpha_{i}$,
(6.1) $\operatorname*{\mathbb{E}}_{S\sim\mu}[p_{i}^{S}\mid i\in S]=1-\prod_{j\in[n]}(1-\alpha_{j}),$
then each agent will obtain a $(1-\prod_{j\in[n]}(1-\alpha_{j}))$-fraction of their ideal utility at equilibrium, and if we have the upper bounds
(6.2) $p_{j}^{S}$ $\leq\frac{1+\alpha_{i}}{2},\,p_{i}^{S}\leq\frac{1+\alpha_{j}}{2}$ $\text{if $S=\{i,j\}$},$
then we obtain $(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2})$-robustness. Lemma˜5.2 showed that allocation probabilities $p_{i}^{S}$ exist that satisfy (6.1) and (6.2) simultaneously.
However, we have not shown how to compute these allocation probabilities. Because there are exponentially many $p_{i}^{S}$’s, we cannot hope to have an efficient algorithm to output all of them at once. Instead, we can only hope for an algorithm that given an $S$, outputs a probability distribution $(p_{i}^{S})_{i\in S}$ in polynomial time that satisfies (6.1) in expectation and (6.2). Then, at a given round in the BRB mechanism, if a set of agents $S^{*}$ bids for the resource, we use the allocation probabilities $(p_{i}^{S^{*}})$ computed from our algorithm run on $S^{*}$.
To compute the allocation probabilities in this fashion, we use the framework presented in [bhalgat2013optimal], originally presented as a framework to design optimal Bayesian Incentive Compatible auctions. This way, we obtain allocation probabilities that always satisfy (6.2), and approximately satisfy (6.1) with high probability.
6.1 Using the Framework of [bhalgat2013optimal] to Compute Allocation Probabilities
Our use of the framework in [bhalgat2013optimal] is similar to the examples they present. For completeness, we show how to use their framework in the context of our problem.
6.1.1 Multiplicative Weights Framework to Solve Linear Programs
The framework of [bhalgat2013optimal] uses the Multiplicative Weights algorithm (MWU) to find feasible solutions to linear programs as done in [arora2012multiplicative]. We first summarize the linear programming environment.
  • A closed set $P\subseteq\mathbb{R}^{m}$ that contains the feasible region.
  • The number of rounds $K$, which will also affect the accuracy of our solution.
  • Matrices $A^{(k)}\in\mathbb{R}^{l\times m}$ and vectors $b^{(k)}\in\mathbb{R}^{l}$ for each $k\in[K]$, each corresponding to a per-round set of linear constraints $A^{(k)}x\geq b^{(k)}$ for $x\in P$.
  • Oracles that solve $\max_{x\in P}y^{\top}A^{(k)}x$ for every given $k\in[K]$ and dual vector $y\geq 0$.
  • A fixed (known) number $\rho$ such that $\|A^{(k)}x-b\|_{\infty}\leq\rho$ for each $k\in[K]$.
The goal is to find solutions $x^{(k)}$ for each $k\in[K]$ such that the total average violation $\frac{1}{K}\sum_{k}(A^{(k)}x^{(k)}-b^{(k)})$ is small (the averaging will approximate the expectation of Eq.˜6.1). To do so, we run MWU for $K$ iterations, maintaining and updating a dual vector $y\in\mathbb{R}^{l}$. Let $y^{(k)}$ be the weight vector at iteration $k$. On iteration $k$, we use our oracle to find $x^{(k)}$ that solves $\max_{x\in P}(y^{(k)})^{\top}A^{(k)}x$. If $(y^{(k)})^{\top}A^{(k)}x^{(k)}<(y^{(k)})^{\top}x^{(k)}$, then we declare that the LP $A^{(k)}x\geq b^{(k)}$ is infeasible. Otherwise, we run a multiplicative weights update on the weight vector $y^{(k)}$ with cost vector $(A^{(k)}x^{(k)}-b^{(k)})/\rho$ to obtain a new weight vector $y^{(k+1)}$, i.e.,
$y^{(k+1)}_{i}=\left(1-\eta\frac{A^{(k)}x^{(k)}-b^{(k)}}{\rho}\right)y^{(k)}_{i}$
This is the same as the multiplicative weights procedure in [arora2012multiplicative] except that the linear programs $A^{(k)}x\geq b^{(k)}$ change with the iteration $k$ as opposed to being fixed throughout. As pointed out by [bhalgat2013optimal], we can still obtain the following guarantee via the same analysis as in [arora2012multiplicative].
Theorem 6.1.
If the algorithm declares that an LP $A^{(k)}x\geq b^{(k)}$ is infeasible, then it is indeed so. Otherwise, let the learning rate used in multiplicative weights be $\eta\in(0,1/2)$. Then,
(6.3) $\frac{1}{K}\sum_{k=1}^{K}\left(A^{(k)}x^{(k)}-b^{(k)}\right)\geq-\rho\left(\eta+\frac{\ln l}{\eta K}\right)$
where the inequality is interpreted coordinate-wise.
6.1.2 Using the Multiplicative Weights Framework to Compute Allocation Probabilities
In this section, we show how to use the setting of subsection˜6.1.1 for finding allocation probabilities for the BRB mechanism. Specifically, we are interested in finding allocation probabilities $(p_{i}^{S})_{S\subseteq[n],i\in S}$ that approximately satisfy (6.1) and (6.2). Equivalently we want
(6.4) $\frac{1}{\alpha_{i}}\sum_{S\subseteq[n]:i\in S}p_{i}^{S}\left(\prod_{j\in S}\alpha_{j}\right)\left(\prod_{j\notin S}(1-\alpha_{j})\right)=1-\prod_{j\in[n]}(1-\alpha_{j})$ $\forall i\in[n],$
and the constraints $\mathcal{F}(S)$ for each $S\subseteq[n]$ to hold, where $\mathcal{F}(S)$ consists of the constraints
$p_{i}^{S}$ $\geq 0$ $\forall i\in S$
$\sum_{i\in S}p_{i}^{S}$ $=1$ if $S\neq\emptyset$
$p_{j}^{S}$ $\leq\frac{1+\alpha_{i}}{2},\,p_{i}^{S}\leq\frac{1+\alpha_{j}}{2}$ $\text{if $S=\{i,j\}$},$
which are linear constraints requiring that $(p_{i}^{S})_{i\in S}$ is indeed a probability distribution over $i\in S$ that satisfy the upper bounds in (6.2). The above constraints define a linear program over variables $p_{i}^{S}$. In terms of the multiplicative weights framework, we encode the allocation probabilities $p_{i}^{S}$ as the vector $x$ and the let the set $P$ be the set of allocation probabilities satisfying $\mathcal{F}(S)$ for each $S\subseteq[n]$.
We now describe an algorithm to compute allocation probabilities for a single round of the BRB mechanism. At a given time, some subset $S^{*}$ of agents bid, and we must produce a probability distribution $(p_{i}^{S^{*}})_{i\in S^{*}}$. At each iteration $k$ of the MWU algorithm, we sample some multiset $\mathcal{S}^{(k)}$ of $C$ i.i.d. samples from [^3] Recall $\mu$ is the distribution of subsets of $[n]$ where each agent $i$ appears with probability $\alpha_{i}$. $\mu$. We let $\mathcal{S}_{i}^{(k)}=\{S\in\mathcal{S}^{(k)}:i\in S\}$ be the sets of $\mathcal{S}^{(k)}$ that agent $i$ is part of. We define an approximate version of (6.4) using a parameter $\delta>0$:
(6.5) $\left|\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}p_{i}^{S}-\left(1-\prod_{j\in[n]}(1-\alpha_{j})\right)\right|\leq\delta$ $\forall i\in[n]$
which form the constraints $A^{(k)}x\geq b^{(k)}$. Specifically, for a fixed agent $i$, we encode the above constraint as the two following (linear) constraints
(6.6) $\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}p_{i}^{S}$ $\geq\max\left\{0,1-\prod_{j\in[n]}(1-\alpha_{j})-\delta\right\}$
(6.7) $-\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}p_{i}^{S}$ $\geq-\min\left\{1,1-\prod_{j\in[n]}(1-\alpha_{j})+\delta\right\}.$
While the maximum with 0 and minimum with $1$ are not conceptually important, they can be used since the $p_{i}^{S}$ are probabilities, and they ensure that $\|A^{(k)}x-b\|_{\infty}\leq 1$ so we can use $\rho=1$ in the multiplicative weights framework.
Let us write down the objective function to the oracle problem, $\max_{x\in P}y^{\top}A^{(k)}x$. Let $y_{i,+}$ be the dual multiplier to (6.6) and let $y_{i,-}$ be the dual multiplier to (6.7). Then, the objective $y^{\top}A^{(k)}x$ is
$\sum_{i\in[n]}\frac{(y_{i,+}-y_{i,-})}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}p_{i}^{S}=\sum_{S\in\mathcal{S}_{i}^{(k)}}\sum_{i\in S}\frac{(y_{i,+}-y_{i,-})p_{i}^{S}}{|\mathcal{S}_{i}^{(k)}|}$
The right-hand side writes the objective as the sum of functions over $S\in\mathcal{S}_{i}^{(k)}$, so the problem decouples. Specifically, to solve the oracle problem $\max_{x\in P}y^{\top}A^{(k)}x$, we just need to solve individual problems $\mathcal{A}^{(k)}(\mathcal{S},y,S)$ for each fixed $S\in\mathcal{S}_{i}^{(k)}$, where $y$ denotes the vector of dual variables $y_{i,+}$ and $y_{i,-}$ and $\mathcal{A}^{(k)}(\mathcal{S},y,S)$ is the following problem.
(6.8) $\mathcal{A}^{(k)}(\mathcal{S},y,S):\quad\max_{(p_{i}^{S})_{i\in S}}\ \sum_{i\in S}\frac{(y_{i,+}-y_{i,-})p_{i}^{S}}{|\mathcal{S}_{i}^{(k)}|}\quad\text{subject to}\quad\mathcal{F}(S)$
We emphasize that (6.8) is simple to solve: it is linear and has at most $n$ variables and $4$ linear constraints.
At the end of the $K$ iterations of MWU, we will either have outputted that some LP was infeasible, or we will have dual vectors $y^{(k)}\in\mathbb{R}^{2n}$ for each $k\in[K]$. In the case that we have outputted that some LP in infeasible, we can output arbitrary $y^{(k)}\in\mathbb{R}^{2n}$. Otherwise, we output $(p_{i}^{S})^{(k)}$ as the solution to each $\mathcal{A}^{(k)}(\mathcal{S}^{(k)},y^{(k)},S)$. Then, when the actual set of bidding agents is $S^{*}$, we allocate according to the probability distribution $\frac{1}{K}\sum_{k=1}^{K}(p_{i}^{S^{*}})^{(k)}$.
We give pseudocode for the above algorithm below. In Algorithm˜1, we describe the procedure to compute the dual vectors. In Algorithm˜2, we describe how to use Algorithm˜1 to compute allocation probabilities.
Algorithm 1 Compute Dual Vectors
0: Fair shares $\alpha_{i}$, number of multiplicative weights iterations $K$, multiplicative weights learning rate $\eta$, number of samples of sets of bidding agents at each iteration $C$, slack parameter $\delta$.
 Initialize dual variables $y_{i,+}^{(1)}=1$ and $y_{i,-}^{(1)}=1$ for each $i\in[n]$.
for $k=1,2,\dots,K$ do
  Sample a multiset $\mathcal{S}^{(k)}$ of $C$ i.i.d. samples from $\mu$.
  For each $S\in\mathcal{S}^{(k)}$, find $(p_{i}^{S})_{i\in S}^{(k)}$ by solving $\mathcal{A}^{(k)}(\mathcal{S}^{(k)},y^{(k)},S)$ defined in (6.8).
  Define $\mathcal{S}_{i}^{(k)}=\{S\in\mathcal{S}^{(k)}:i\in S\}$.
  Define the costs $c_{i,+}^{(k)}=\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}p_{i}^{S}-\max\left\{0,1-\prod_{j\in[n]}(1-\alpha_{j})-\delta\right\}$ and $c_{i,-}^{(k)}=-\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}p_{i}^{S}+\min\left\{1,1-\prod_{j\in[n]}(1-\alpha_{j})+\delta\right\}$.
  Update the dual variables as as in the MWU algorithm: $y_{i,+}^{(k+1)}=(1-\eta c_{i,+})y_{i,+}^{(k)}$ and $y_{i,-}^{(k+1)}=(1-\eta c_{i,-})y_{i,-}^{(k)}$.
end for
 Dual vectors $y^{(k)}=(y_{i,+}^{(k)},y_{i,-}^{(k)})_{i\in[n]}$ for each $k\in K$ and samples $\mathcal{S}^{(k)}$.
Algorithm 2 Compute Allocation Probabilities
0: Set $S^{*}$ of bidding agents.
 Obtain dual vectors $y^{(k)}$ and samples $\mathcal{S}^{(k)}$ for each $k\in[K]$ from Algorithm˜1.
 Obtain $(p_{i}^{S^{*}})^{(k)}$ by solving $\mathcal{A}^{(k)}(\mathcal{S}^{(k)},y^{(k)},S^{*})$ for each $k\in[K]$.
 Output $(p_{i}^{S^{*}}):=\frac{1}{K}\sum_{k=1}^{K}(p_{i}^{S^{*}})^{(k)}$.
The following lemma, which we prove in subsection˜E.9, shows that the allocation probabilities approximately induce the correct interim allocation probabilities with high probability, that holds independently for each requesting set $S^{*}$.
Lemma 6.2.
Suppose we run Algorithm˜2 with $S^{*}\sim\mu$. When we run Algorithm˜1 as a subroutine, given a fixed parameter $K$, choosing the other parameters as
(6.9) $\eta=\min\left\{\frac{1}{2},\sqrt{\frac{\ln n}{K}}\right\},C=\left\lceil K\left(\frac{\ln nK}{\underline{\alpha}}\right)^{2}\right\rceil,\delta=\sqrt{\frac{3\ln nK}{\underline{\alpha}C}}$
there is an event $E$ contained in the $\sigma$-algebra generated by the samples $(\mathcal{S}^{(k)})_{k\in[K]}$ of probability at least $1-O\left(\frac{1}{n^{2}K^{2}}\right)$ such that on $E$, for every agent $i$,
$\left|\operatorname*{\mathbb{E}}_{\begin{subarray}{c}S^{*}\sim\mu\end{subarray}}\left[p_{i}^{S^{*}}\,\middle|\,i\in S^{*},(\mathcal{S}^{(k)})_{k\in[K]}\right]-\left(1-\prod_{j\in[n]}(1-\alpha_{j})\right)\right|\leq O\left(\sqrt{\frac{\log n}{K}}\right)$
where $\underline{\alpha}=\min_{j}\alpha_{j}$.
Also, when choosing such parameters the algorithm runs in time $O(\operatorname{poly}(K,n,1/\underline{\alpha}))$.
6.2 BRB with Online Computation of Allocation Probabilities
We now describe how to run our BRB mechanism while efficiently computing the allocation probabilities online. At the beginning of the mechanism, we use Algorithm˜1 to obtain dual vectors $(y^{(k)})$. Then, at each time step $t$, given the set of bidding agents $S^{t}$, we can use the dual vectors $(y^{(k)})$ to obtain allocation probabilities $(p_{i}^{S^{t}})$ to be used in round $t$ using Algorithm˜2. The formal description of our algorithm is in Algorithm˜3.
Our argument to show that this maintains the robustness and equilibrium guarantees goes as follows. The allocation probabilities $p_{i}^{S^{*}}$ outputted by Algorithm˜2 necessarily satisfy $\mathcal{F}(S^{*})$ so they will satisfy (6.2). We will use Lemma˜6.2 to show that (6.1) is approximately satisfied with high probability. The equilibrium utility and robustness will then follow.
Algorithm 3 BRB with Efficient Computation of Allocation Probabilities $p_{i}^{S}$.
 Obtain dual vectors $y^{(k)}$ and samples $\mathcal{S}^{(k)}$ for each $k\in[K]$ from Algorithm˜1.
 We run BRB as usual and compute the allocation probabilities online from $y^{(k)}$ and $\mathcal{S}^{(k)}$ as follows.
for $t=1,2,\dots,T$ do
  Observe the set of bidding agents $S^{t}$.
  Sample $k^{t}\sim[K]$.
  Use the allocation probabilities $(p_{i}^{S^{t}})^{(k^{t})}_{i\in S^{t}}$ given by solving $\mathcal{A}^{(k^{t})}(\mathcal{S}^{(k^{t})},y^{(k^{t})},S^{t})$.
end for
Formally, we have the following, which we prove in subsection˜E.9. Due to the error by the sampling being proportional to $\tilde{O}(1/\sqrt{T})$, we get the same performance as in Theorem˜3.6.
Theorem 6.3.
 [^4] Outside of this section, we use big $O$ notation to only indicate dependence on $T$. In all our uses, this suppresses only polynomial dependence on $n$ and $(1/\alpha_{j})_{j\in[n]}$. In this section, we explicitly keep the $\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})$ factors in the big $O$ notation.
Algorithm˜3 with $K=T$ and other parameters set as in Lemma˜6.2 has the following guarantees. Each player $i$ playing an $\alpha_{i}$-aggressive strategy is an $O\left(\sqrt{\frac{\log T}{T}}\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})\right)$-equilibrium in which player $i$ receives utility
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\left(1-\prod_{j\in[n]}(1-\alpha_{j})\right)v_{i}^{\star}-O\left(\sqrt{\frac{\log T}{T}}\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})\right)$
with probability at least $1-O(\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})/T^{2})$. Furthermore, regardless of the behavior of other agents $j\neq i$, if agent $i$ plays an $\alpha_{i}$-aggressive strategy, she is guaranteed utility
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2}-O\left(\sqrt{\frac{\log T}{T}}\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})\right)$
with probability at least $1-O(\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})/T^{2})$.
A Warm Up: Repeated Allocation with Two Agents
To motivate the design of our mechanism and show how we can achieve good robust equilibria as outlined in subsection˜2.1, we first give a simplified version of our mechanism in the special case where there are only two agents. Our mechanism lets each agent decide to bid for the resource or not in each round, up to some limited number of bids. It makes sense that when only one agent bids (and has not exceeded her bid budget), she is guaranteed the item. What is unclear is who gets allocated when both agents bid. We show that carefully randomizing this allocation leads to the optimal guarantee of Lemma˜2.2, both in terms of the equilibrium and the robust setting, i.e., $\lambda_{\mathtt{NASH}}=\lambda_{\mathtt{ROB}}=1-(1-\alpha_{1})(1-\alpha_{2})$. To extend this result to $n$ agents, we need to set exponentially many parameters (the allocations when any subset $S$ agents bids), which we do by strengthening Border’s theorem in section˜4.
Our mechanism proceeds as follows: Each agent $i$ starts with $\alpha_{i}T$ bid tokens. At each round, each agent can bid 0 or $1$. If both agents bid 0, no one receives the item. If only one agent bids $1$, then she receives the item in that round. If both agents bid $1$, then we allocate the item to agent $1$ with probability $p_{1}^{\{1,2\}}=p$ and to agent $2$ with probability $p_{2}^{\{1,2\}}=1-p$. Regardless of who wins the item, each agent $i$ pays their bid. If an agent’s budget becomes 0, they can not bid.
Proposition A.1 (Informal).
Proof 4 (Proof sketch).
First, observe that the expected spending of each agent $i$ under this strategy profile is indeed at most $B_{i}[1]=\alpha_{i}T$, since each agent will be bidding i.i.d. $\mathrm{Bernoulli}(\alpha_{i})$. Next, note that there is nothing that agent $1$ can do to affect player $2$’s behavior. She must solve her own stochastic control problem where her policy does not affect agent $2$’s behavior. At any time $t$, if agent $1$ bids, her probability of winning is $p_{1}=\alpha+(1-\alpha)p$, where the $\alpha$ term is the probability that agent $2$ does not bid, and the $(1-\alpha)p$ term is the probability that agent $2$ bid, and agent $1$ wins give both agents bid. Thus, agent $1$’s control problem reduces to selecting $\alpha_{1}T$ time periods in which to bid, and she wins a $p_{1}$ fraction of them in expectation. Agent $1$ should bid on her highest-valued $\alpha_{i}T$ time periods, which is precisely what bidding whenever her value $V_{i}[t]$ is in the top $\alpha_{i}$-quantile of her distribution does (in expectation).
where $p_{1}=\alpha+(1-\alpha)p$ is the probability agent $1$ wins a round conditioned on her bidding. Symmetrically, player $2$ gets expected utility $\frac{1}{T}\sum_{t}\mathbb{E}[U_{2}[t]]=p_{2}v_{2}^{\star}$ where $p_{2}=(1-\alpha)+\alpha(1-p)=1-\alpha p$. Setting $p=1-\alpha$ guarantees both agents $1-\alpha(1-\alpha)$ fraction of their ideal utility. This achieves $\lambda_{\mathtt{NASH}}=1-\alpha(1-\alpha)$ and by Lemma˜2.2, this is the best factor any allocation rule can get.
The choice of $p=1-\alpha$ might seem unusual, especially when $\alpha$ is close to 0 or $1$ where one agent wins most times. The intuition behind this is as follows: if agent $1$ has small fair share and agent $2$ has a high one, then using a high $p$ and favoring agent $1$ has a small effect on agent $2$, most of whose bids are uncontested. Therefore, if we want equal outcomes, i.e., $p_{1}=p_{2}$, we have to favor the agent with the small fair share, breaking most ties in her favor.
Proposition A.2 (Informal).
This means we achieve $\lambda_{\mathtt{ROB}}=1-(1-\alpha_{1})(1-\alpha_{2})$ which also matches the upper bound of any allocation of Lemma˜2.2.
In coming sections, we generalize this mechanism for $n\geq 2$ players, where the main difficulty is resolving concurrent bids for any subset $S\subseteq[n]$ with $|S|\geq 2$. Since there are exponentially many such subsets, it is much harder to find a good or simple allocation rule that has the same properties as our choice of $p$ above. We will show that we can get the same optimal factor under equilibrium performance, $\lambda_{\mathtt{NASH}}=1-\prod_{i}(1-\alpha_{i})$, using Border’s theorem. For the robust guarantee, we show that a simple application of Border’s theorem is not enough. Instead, we have to consider more restrictions on how we handle concurrent bids, which leads to the strengthening of Border’s theorem in Lemma˜5.2. Unfortunately, the robust guarantee weakens for $n>2$, and we get $\lambda_{\mathtt{ROB}}\approx 1/2$, which we show is tight under any rule for resolving concurrent bids.
B Better Equilibrium Guarantees